OpenStudy (cruxim):

Write the equation of the line perpendicular to the line x - 3y = 9 and passing through the point (3,5).

OpenStudy (anonymous):

@Cruxim First, change the equation to slope intercept form: \[y=mx+b\] Second, follow the point slope form with the set of points given. From the first step, you will get the slope of the line. Replace it in this equation below: \[y-y _{1}=m(x-x _{1})\] Tell me what answer did you get. :)

OpenStudy (anonymous):

By the way, welcome to Open Study! :)

OpenStudy (cruxim):

Ah thank you for the welcome! For slope intercept I got: y=x/3-3 For x I got -12 and for y I got -4